# Quantum many-body analysis of a spin-2 bosons with two-body inelastic decay

### v1

@@ III GENERAL FORMS OF STEADY STATES
where $2N\leq F_{z}\leq 2N$. In the following, we write the density matrix in the $N$-particle subspace as $\hat{\rho}_{N}$ and express the density matrix for the total Hilbert space as
$\hat{\rho}=\oplus_{N=0}^{N_{\textrm{ini}}}\hat{\rho}_{N},$ (18)
where $N_{\textrm{ini}}$ is the maximum number of atoms contained in the initial state. We will prove that $\hat{\rho}_{N}$ has the form
$\hat{\rho}_{N}=\sum_{F_{z},F_{z}^{\prime}}c_{F_{z}F_{z}^{\prime}}|F=2N,F_{z}\rangle\langle F=2N,F_{z}^{\prime}|,$ (19)
where $c_{F_{z}F_{z}^{\prime}}$ is a positive semidefinite Hermitian matrix.
To begin with, we prove that Eq. (19) is a sufficient condition for Eq. (16). The states in Eq. (17) are expressed as [22]
$|F=2N,F_{z}\rangle=Z^{-1/2}(\hat{F}_{-})^{\Delta F}(\hat{a}^{\dagger}_{2})^{N}|vac\rangle,$ (20)
where $\hat{F}_{-}=2(\hat{a}^{\dagger}_{1}\hat{a}_{2}+\hat{a}_{-2}^{\dagger}\hat{a}_{-1})+\sqrt{6}(\hat{a}^{\dagger}_{0}\hat{a}_{1}+\hat{a}^{\dagger}_{-1}\hat{a}_{0})$, $\Delta F=2N-F_{z}$, $Z=N!(4N)!\Delta F!/(4N-\Delta F)!$ is a normalization constant, and $|vac\rangle$ is the vacuum state. From the form of $\hat{A}_{00}=(\hat{a}_{0}\hat{a}_{0}-2\hat{a}_{1}\hat{a}_{-1}+2\hat{a}_{2}\hat{a}_{-2})/\sqrt{5}$, we find $\hat{A}_{00}(\hat{a}^{\dagger}_{2})^{N}|vac\rangle=0$ and $\hat{A}_{00},\hat{F}_{-}]=0$, which gives $\hat{A}_{00}(\hat{F}_{-})^{\Delta F}(\hat{a}^{\dagger}_{2})^{N}|vac\rangle=0$. Similarly, $\hat{A}_{2\mathcal{M}}$ satisfies $\hat{A}_{2\mathcal{M}}(\hat{a}^{\dagger}_{2})^{N}|vac\rangle=0$. From the commutation relations between $\hat{A}_{2\mathcal{M}}$ and $\hat{F}_{-}$,
@@
we obtain $\hat{A}_{2\mathcal{M}}(\hat{F}_{-})^{\Delta F}(\hat{a}^{\dagger}_{2})^{N}|vac\rangle=0$. Thus, we have shown that
$\hat{A}_{\mathcal{FM}}|F=2N,F_{z}\rangle=0,$ (22)
for $\mathcal{F}=0$ and 2, and therefore, Eq. (19) satisfies Eq. (16).
Next, we show that Eq. (19) is a necessary condition for Eq. (16). First, we prove the lemma that Eq. (16) leads to
$\hat{A}_{\mathcal{FM}}\hat{\rho}_{N}\hat{A}_{\mathcal{FM}}^{\dagger}=0$ (23)
for $\mathcal{F}=0,2$ and for all possible $N$. We start from the dissipation from the $N=2$ space to the $N=0$ space. From $\hat{A}_{\mathcal{FM}}^{\dagger}\hat{A}_{\mathcal{FM}}\rho_{0}=\rho_{0}\hat{A}_{\mathcal{FM}}^{\dagger}\hat{A}_{\mathcal{FM}}=0$, Eq. (16) gives
$\sum_{\mathcal{F}=0,2}\frac{b_{\mathcal{F}}}{4}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}2\hat{A}_{\mathcal{FM}}\hat{\rho}_{2}\hat{A}_{\mathcal{FM}}^{\dagger}=0.$ (24)
Taking the trace of both sides of Eq. (24), we obtain
$\textrm{Tr}\left[\sum_{\mathcal{F}=0,2}\frac{b_{\mathcal{F}}}{4}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}2\hat{A}_{\mathcal{FM}}\hat{\rho}_{2}\hat{A}_{\mathcal{FM}}^{\dagger}\right]$ (25)
$=\sum_{\mathcal{F}=0,2}\frac{b_{\mathcal{F}}}{4}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}2\textrm{Tr}\left[\hat{A}_{\mathcal{FM}}^{\dagger}\hat{A}_{\mathcal{FM}}\hat{\rho}_{2}\right]$
@@
From the positive definiteness of $\hat{A}_{\mathcal{FM}}^{\dagger}\hat{A}_{\mathcal{FM}}$, each term of the sum in Eq (LABEL:tracing) must vanish; therefore, Eq. (23) is satisfied for $N=2$. For the dissipation from $N=4$ to $N=2$,
$\sum_{\mathcal{F}=0,2}\frac{b_{\mathcal{F}}}{4}$ $\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}\left(2\hat{A}_{\mathcal{FM}}\hat{\rho}_{4}\hat{A}_{\mathcal{FM}}^{\dagger}\right$. (26)
- $\left.\hat{A}_{\mathcal{FM}}^{\dagger}\hat{A}_{\mathcal{FM}}\hat{\rho}_{2}-\hat{\rho}_{2}\hat{A}_{\mathcal{FM}}^{\dagger}\hat{A}_{\mathcal{FM}}\right)=0$
holds from Eq. (16), where the second and third terms in the bracket vanish according to Eq. (LABEL:tracing). Therefore, by a similar procedure, we can show that Eq. (23) also holds for $N=4$. By mathematical induction, Eq. (23) holds for $N=2,4,...,N_{\textrm{ini}}$. Also, for odd values of $N$, we can similarly show that Eq. (23) holds. Thus, when the density matrix $\hat{\rho}_{N}$ satisfies Eq. (16), Eq. (23) holds for all possible $N$.
In general, the density matrix $\hat{\rho}_{N}$ is written as a mixture of pure states $\{|\psi_{N}^{(k)}\rangle\}_{k=1,2,...,K}$ as
$\hat{\rho}_{N}=\sum_{k=1}^{K}P_{N}^{(k)}|\psi_{N}^{(k)}\rangle\langle\psi_{N}^{(k)}|,$ (27)
where $P_{N}^{(k)}$ is positive. When the density matrix satisfying Eq. (23) is expressed by Eq. (27),
$\hat{A}_{\mathcal{FM}}|\psi_{N}^{(k)}\rangle$ $=0~~~~(\mathcal{F}=0,2),$ (28)
must hold for each pure state $|\psi_{N}^{(k)}\rangle$. The states in Eq. (17) satisfy Eq. (28) as shown in Eq. (22). We can show that no other states satisfy Eq. (28) as follows. Equation (28) leads to
$\sum_{\mathcal{F}=0,2}\frac{g_{\mathcal{F}}}{2V^{\textrm{eff}}}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}\hat{\mathcal{A}}^{\dagger}_{\mathcal{FM}}\hat{\mathcal{A}}_{\mathcal{FM}}|\psi_{N}^{(k)}\rangle=0.$ (29)
@@
$\lambda=$ $\frac{g_{2}}{7}[2N(2N+1)-F(F+1)]$ (30)
$+$ $\frac{7g_{0}-10g_{2}}{7}N_{S}(N+N_{0}+3)$
from Eq. (14). The quantum numbers $N_{S}$ and $F$ that give $\lambda=0$ are only $N_{S}=0$ and $F=2N$; that is, Eq. (29) is satisfied only by the states in Eq. (17). Therefore, Eq. (28) is satisfied only by the states in Eq. (17). Thus, Eq. (16) is satisfied if and only if the state has the form of Eq. (19). In conclusion, the steady state that satisfies Eqs. (11) and (16) for $q=0$ must have the form of Eq. (18) with Eq. (19).
Finally, we consider the steady state for $q\neq 0$. Since Eq. (16) is independent of the value of $q$, the steady-state density matrix $\hat{\rho}$ must be a mixture of the states given by Eq. (17). However, because of the quadratic Zeeman term in the Hamiltonian,
$\hat{Q}=q\sum_{m=-2}^{2}m^{2}\hat{a}^{\dagger}_{m}\hat{a}_{m},$ (31)
some of the states in Eq. (17) are not the eigenstates of the Hamiltonian; for these states, Eq. (13) does not hold. Among the states in Eq. (17), only the states
$|F=2N,F_{z}$ $=-2N\rangle$ $=\frac{(\hat{a}^{\dagger}_{-2})^{N}}{\sqrt{N!}}|vac\rangle,$ (32)
$|F=2N,F_{z}$ $=2N\rangle$ $=\frac{(\hat{a}^{\dagger}_{2})^{N}}{\sqrt{N!}}|vac\rangle,$
@@
$|F=2N,F_{z}$ $=-2N+1\rangle$ $=\frac{\hat{a}^{\dagger}_{-1}(\hat{a}^{\dagger}_{-2})^{N-1}}{\sqrt{(N-1)!}}|vac\rangle,$ (33)
$|F=2N,F_{z}$ $=2N-1\rangle$ $=\frac{\hat{a}^{\dagger}_{1}(\hat{a}^{\dagger}_{2})^{N-1}}{\sqrt{(N-1)!}}|vac\rangle,$
are eigenstates of the Hamiltonian with eigenvalues $E_{q=0}+4qN$ for Eq. (32) and $E_{q=0}+4q(N-1)+q$ for Eq. (33), where $E_{q=0}$ is the eigenvalue for $q=0$. Thus, if the density matrix in Eq. (19) is constructed from the states in Eqs. (32) and (33) such that Eq. (13) is satisfied, we obtain a steady state for $q\neq 0$.
@@ IV NUMERICAL RESULTS
Table 1: Interaction and dissipation parameters $g_{\mathcal{F}}$ and $b_{\mathcal{F}}$ of $^{87}\textrm{Rb}$, taken from Refs. [13] and [20], respectively.
Interaction $10^{-11}\textrm{cm}^{3}/\textrm{s}$ Loss $10^{-11}\textrm{cm}^{3}/\textrm{s}$
@@
We study the time evolution of the system by numerically solving the master equation (4) using the fourth-order Runge-Kutta method. The initial state is a Fock state $(\hat{a}^{\dagger}_{0})^{N_{\textrm{ini}}}|\textrm{vac}\rangle/\sqrt{N_{\textrm{ini}}!}$ occupying the $m=0$ state, as in the experiment in Ref. [21]. Here, we restrict ourselves to $N_{\textrm{ini}}=20$, which requires $\sim 10^{9}$ matrix elements to represent the density matrix. We consider an atomic gas of $^{87}\mathrm{Rb}$ confined in an isotropic harmonic potential with a trap frequency $\omega/(2\pi)=2~\mathrm{kHz}$. We use the single-particle ground-state wave function for the single-mode approximation, and the effective volume is given by $V^{\textrm{eff}}\simeq 2.2\times 10^{-13}~\mathrm{cm}^{3}$. The interaction and dissipation parameters are set to the experimentally measured values given in Table 1. For these interaction parameters, the ground state of the Hamiltonian for $q=0$ is the state where all atoms form singlet pairs.
As the system evolves over time, the total density matrix for the system has the form $\hat{\rho}=\oplus_{N=0}^{20}\hat{\rho}_{N}$, where $N$ are even integers. The expectation value of an operator $\hat{A}$ is given by $\langle\hat{A}\rangle=\textrm{Tr}[\hat{A}\hat{\rho}]/\textrm{Tr}[\hat{\rho}$], and that of $\hat{A}$ within the $N$-particle subspace is defined by $\langle\hat{A}\rangle_{N}=\textrm{Tr}[\hat{A}\hat{\rho}_{N}]/\textrm{Tr}[\hat{\rho}_{N}]$. The spin operator in the $j$-direction ($j=x,y,z$) is expressed as $\hat{F}_{j}=\sum_{m,m^{\prime}=-2}^{2}[f_{j}]_{m,m^{\prime}}\hat{a}_{m}^{\dagger}\hat{a}_{m^{\prime}}$, where $f_{j}$ are spin-2 matrices. For the present initial state, $\langle\hat{F}_{j}\rangle$ and $\langle\hat{F}_{j}\rangle_{N}$ always vanish as the system evolves. We define the transverse magnetization as $S_{\perp}=\sqrt{\langle\hat{F}_{x}^{2}+\hat{F}_{y}^{2}\rangle}$. To examine whether the state converges to the fully magnetized steady state in Eq. (18) with Eq. (17), we introduce the normalized spin-squared operator as
$\hat{\textbf{F}}^{2}_{\textrm{norm}}=\frac{\hat{F}_{x}^{2}+\hat{F}_{y}^{2}+\hat{F}_{z}^{2}}{2\hat{N}(2\hat{N}+1)}.$ (34)
To avoid divergence arising from $N=0$, we define the density matrix restricted to the $N>0$ subspace as $\hat{\rho}_{N>0}=\oplus_{N=2}^{20}\hat{\rho}_{N}$. The expectation value of Eq. (34) is taken with respect to this restricted state as
$F^{2}_{\textrm{norm}}=\frac{\textrm{Tr}[\hat{\textbf{F}}^{2}_{\textrm{norm}}\hat{\rho}_{N>0}]}{\textrm{Tr}[\hat{\rho}_{N>0}]}.$ (35)
This quantity becomes unity for the state in Eq. (18) with Eq. (17). Let $P_{N}$ be the probability of finding the system in a state with $N$ particles,
$P_{N}=\frac{\textrm{Tr}[\hat{\rho}_{N}]}{\textrm{Tr}[\hat{\rho}]}$. (36)
@@
$Q_{m}^{(N)}=\frac{\textrm{Tr}[\hat{\mathcal{P}}_{m}\hat{\rho}_{N}]}{\textrm{Tr}[\hat{\rho}_{N}]},$ (37)
@@
Figure 1: Time evolution for $q=0$. (a) Average particle number $\langle\hat{N}\rangle$ and normalized total magnetization $F^{2}_{\textrm{norm}}$. (b) Particle number distribusions $P_{N}$ at $t=0,0.132$, and $1.76~\mathrm{s}$. (c) Distribusions $Q_{m}^{(N)}$ of the $z$-magnetization at $t=0.132$ and $1.76~\mathrm{s}$. The upper and lower panels show the distributions in the subspaces $N=6$ and $N=16$, respectively. The dotted lines correspond to Eq. (38).

### v2

@@ III.2 General forms of steady states
where $2N\leq F_{z}\leq 2N$. We define a density operator that only consists of the states with maximum total spin as
$\hat{\rho}^{\mathrm{ss}}_{N}\equiv\sum_{F_{z},F_{z}^{\prime}}c_{F_{z}F_{z}^{\prime}}|F=2N,F_{z}\rangle\langle F=2N,F_{z}^{\prime}|,$ (25)
where $c_{F_{z}F_{z}^{\prime}}$ is an element of a positive semidefinite Hermitian matrix. Using Eq. (25), we define the density operator as
$\hat{\rho}^{\mathrm{ss}}\equiv\oplus_{N=0}^{N_{\mathrm{ini}}}\hat{\rho}^{\mathrm{ss}}_{N}.$ (26)
We will prove that Eq. (26) is a necessary and sufficient condition for a steady-state solution of Eq. (10) satisfying
$\frac{1}{i\hbar}$ $\left[\hat{H},\hat{\rho}^{\mathrm{ss}}\right]+\sum_{\mathcal{F}=0,2}\frac{b_{\mathcal{F}}}{4V^{\mathrm{eff}}}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}\left(2\hat{A}_{\mathcal{FM}}\hat{\rho}^{\mathrm{ss}}\hat{A}_{\mathcal{FM}}^{\dagger}\right$. (27)
$\left.-\hat{A}_{\mathcal{FM}}^{\dagger}\hat{A}_{\mathcal{FM}}\hat{\rho}^{\mathrm{ss}}-\hat{\rho}^{\mathrm{ss}}\hat{A}_{\mathcal{FM}}^{\dagger}\hat{A}_{\mathcal{FM}}\right)=0.$
First, we show that Eq. (26) is a sufficient condition for Eq. (27). The commutator in Eq. (27) vanishes,
$\frac{1}{i\hbar}\left[\hat{H},\hat{\rho}^{\mathrm{ss}}\right]=0,$ (28)
since the bases contained in Eq. (25) are eigenstates of $\hat{H}$ with the same eigenvalues. The state in Eq. (24) is expressed as [22]
$|F=2N,F_{z}\rangle=Z^{-1/2}(\hat{F}_{-})^{\Delta F}(\hat{a}^{\dagger}_{2})^{N}|\mathrm{vac}\rangle,$ (29)
where $\hat{F}_{-}=2(\hat{a}^{\dagger}_{1}\hat{a}_{2}+\hat{a}_{-2}^{\dagger}\hat{a}_{-1})+\sqrt{6}(\hat{a}^{\dagger}_{0}\hat{a}_{1}+\hat{a}^{\dagger}_{-1}\hat{a}_{0})$, $\Delta F=2N-F_{z}$, and $Z=N!(4N)!\Delta F!/(4N-\Delta F)!$. From the form of $\hat{A}_{00}=(\hat{a}_{0}\hat{a}_{0}-2\hat{a}_{1}\hat{a}_{-1}+2\hat{a}_{2}\hat{a}_{-2})/\sqrt{5}$, we find $\hat{A}_{00}(\hat{a}^{\dagger}_{2})^{N}|\mathrm{vac}\rangle=0$ and $\hat{A}_{00},\hat{F}_{-}]=0$, which gives $\hat{A}_{00}(\hat{F}_{-})^{\Delta F}(\hat{a}^{\dagger}_{2})^{N}|\mathrm{vac}\rangle=0$. Similarly, $\hat{A}_{2\mathcal{M}}$ satisfies $\hat{A}_{2\mathcal{M}}(\hat{a}^{\dagger}_{2})^{N}|\mathrm{vac}\rangle=0$. From the commutation relations between $\hat{A}_{2\mathcal{M}}$ and $\hat{F}_{-}$,
@@
we obtain $\hat{A}_{2\mathcal{M}}(\hat{F}_{-})^{\Delta F}(\hat{a}^{\dagger}_{2})^{N}|\mathrm{vac}\rangle=0$. Thus, we have shown that
$\hat{A}_{\mathcal{FM}}|F=2N,F_{z}\rangle=0$ (31)
for $\mathcal{F}=0$ and $\mathcal{F}=2$, clearly indicating that the second term on the left-hand side of Eq. (27) also vanishes,
$\sum_{\mathcal{F}=0,2}\frac{b_{\mathcal{F}}}{4V^{\mathrm{eff}}}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}\left(2\hat{A}_{\mathcal{FM}}\hat{\rho}^{\mathrm{ss}}\hat{A}_{\mathcal{FM}}^{\dagger}\right$. (32)
$\left.-\hat{A}_{\mathcal{FM}}^{\dagger}\hat{A}_{\mathcal{FM}}\hat{\rho}^{\mathrm{ss}}-\hat{\rho}^{\mathrm{ss}}\hat{A}_{\mathcal{FM}}^{\dagger}\hat{A}_{\mathcal{FM}}\right)=0.$
Next, we prove that Eq. (26) is a necessary condition for Eq. (27). In the steady state satisfying Eq. (13), the time derivative of the average particle number $d\langle\hat{N}\rangle/dt$ must be zero. Using the master equation (10) and the commutation relation
$\hat{N},\hat{A}_{\mathcal{FM}}]=-2\hat{A}_{\mathcal{FM}}$, (33)
we can express the time derivative of the average particle number as
$\frac{d\langle\hat{N}\rangle}{dt}$ $=\mathrm{Tr}\left[\hat{N}\frac{d\hat{\rho}(t)}{dt}\right]$ (34)
$=-\mathrm{Tr}\left[\sum_{\mathcal{F}=0,2}\frac{b_{\mathcal{F}}}{V^{\mathrm{eff}}}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}\hat{A}_{\mathcal{FM}}\hat{\rho}\hat{A}_{\mathcal{FM}}^{\dagger}\right]$
$=-\sum_{\mathcal{F}=0,2}\frac{b_{\mathcal{F}}}{V^{\mathrm{eff}}}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}\mathrm{Tr}\left[\hat{A}_{\mathcal{FM}}^{\dagger}\hat{A}_{\mathcal{FM}}\hat{\rho}\right$]
@@
From the positive definiteness of $\hat{A}_{\mathcal{FM}}^{\dagger}\hat{A}_{\mathcal{FM}}$, each term of the sum in the third line of Eq. (34) must vanish. Equation (34) therefore leads to
$\hat{A}_{\mathcal{FM}}\hat{\rho}\hat{A}_{\mathcal{FM}}^{\dagger}=0,$ (35)
for $\mathcal{F}=0,2$.
In general, the density operator $\hat{\rho}$ is written as a mixture of pure states $\{|\psi_{k}\rangle\}_{k=1,2,\dots,K}$ as
$\hat{\rho}=\sum_{k=1}^{K}P_{k}|\psi_{k}\rangle\langle\psi_{k}|,$ (36)
where $P_{k}$ is positive. When the density operator satisfying Eq. (35) is expressed by Eq. (36),
$\hat{A}_{\mathcal{FM}}|\psi_{k}\rangle$ $=0~~~~(\mathcal{F}=0,2)$ (37)
must hold for each pure state $|\psi_{k}\rangle$. The states in Eq. (24) satisfy Eq. (37), as shown in Eq. (31). We can show that no other states satisfy Eq. (37) as follows. Equation (37) leads to
$\sum_{\mathcal{F}=0,2}\frac{b_{\mathcal{F}}}{2V^{\mathrm{eff}}}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}\hat{A}^{\dagger}_{\mathcal{FM}}\hat{A}_{\mathcal{FM}}|\psi_{k}\rangle=0.$ (38)
@@
$\lambda=$ $\frac{b_{2}}{14V^{\mathrm{eff}}}[2N(2N+1)-F(F+1)]$ (39)
$+$ $\frac{7b_{0}-10b_{2}}{35V^{\mathrm{eff}}}N_{S}(N+N_{0}+3),$
from Eq. (11). The quantum numbers $N_{S}$ and $F$ that give $\lambda=0$ are only $N_{S}=0$ and $F=2N$; that is, Eq. (38) is satisfied only by superpositions of the states in Eq. (24). Therefore, Eq. (37) is satisfied only by the states in Eq. (24). Thus, Eq. (26) is a necessary condition for Eq. (27). In conclusion, Eq. (13) for $q=0$ is satisfied if and only if the density operator has the form of Eqs. (25) and (26).
Finally, we consider the steady state for $q\neq 0$. The above proof is valid also for $q\neq 0$ if all of the bases used in Eq. (25) are eigenstates of $\hat{H}$ with the same eigenvalue. However, because of the quadratic Zeeman term in the Hamiltonian, some of the states in Eq. (24) are not eigenstates of the Hamiltonian; for these states, Eq. (28) does not hold. Among the states in Eq. (24), only the states
$|F=2N,F_{z}$ $=-2N\rangle$ $=\frac{(\hat{a}^{\dagger}_{-2})^{N}}{\sqrt{N!}}|\mathrm{vac}\rangle,$ (40)
$|F=2N,F_{z}$ $=2N\rangle$ $=\frac{(\hat{a}^{\dagger}_{2})^{N}}{\sqrt{N!}}|\mathrm{vac}\rangle,$
@@
$|F=2N,F_{z}$ $=-2N+1\rangle$ $=\frac{\hat{a}^{\dagger}_{-1}(\hat{a}^{\dagger}_{-2})^{N-1}}{\sqrt{(N-1)!}}|\mathrm{vac}\rangle,$ (41)
$|F=2N,F_{z}$ $=2N-1\rangle$ $=\frac{\hat{a}^{\dagger}_{1}(\hat{a}^{\dagger}_{2})^{N-1}}{\sqrt{(N-1)!}}|\mathrm{vac}\rangle,$
are eigenstates of the Hamiltonian with eigenvalues $E_{q=0}+4qN$ for Eq. (40) and $E_{q=0}+4q(N-1)+q$ for Eq. (41), where $E_{q=0}$ is the eigenvalue of the Hamiltonian for $q=0$. Thus, if the density operator in Eq. (25) is constructed from the states in Eqs. (40) and (41) such that Eq. (28) is satisfied, we obtain a steady state for $q\neq 0$.
@@ IV NUMERICAL RESULTS
Table 1: Interaction and dissipation parameters $g_{\mathcal{F}}$ and $b_{\mathcal{F}}$ of $^{87}\mathrm{Rb}$, taken from Refs. [13] and [20], respectively.
Interaction $10^{-11}\mathrm{cm}^{3}/\mathrm{s}$ Loss $10^{-11}\mathrm{cm}^{3}/\mathrm{s}$
@@
We study the time evolution of the system by numerically solving the master equation (4) using the fourth-order Runge-Kutta method. The initial state is a Fock state
$|\Psi_{\mathrm{ini}}\rangle=\frac{(\hat{a}^{\dagger}_{0})^{N_{\mathrm{ini}}}}{\sqrt{N_{\mathrm{ini}}!}}|\mathrm{vac}\rangle,$ (42)
occupying the $m=0$ state, as in the experiment in Ref. [21]. Here, we restrict ourselves to $N_{\mathrm{ini}}=20$, which requires $\sim 10^{9}$ matrix elements to represent the density operator. We consider an atomic gas of $^{87}\mathrm{Rb}$ confined in an isotropic harmonic potential with a trap frequency $\omega/(2\pi)=2~\mathrm{kHz}$. We use the single-particle ground-state wave function for the single-mode approximation, and the effective volume is given by $V^{\rm eff}=\left[\int|\psi_{\mathrm{SMA}}(\bm{r})|^{4}d\bm{r}\right]^{-1}\simeq 2.2\times 10^{-13}~\mathrm{cm}^{3}$. The interaction and dissipation parameters are set to the experimentally measured values given in Table 1. For these interaction parameters, the ground state of the Hamiltonian for $q=0$ is the state where all atoms form singlet pairs [22].
As the system evolves over time, the total density operator for the system has the form $\hat{\rho}=\oplus_{N=0}^{20}\hat{\rho}_{N}$, where $N$ are even integers. The expectation value of an operator $\hat{A}$ is given by $\langle\hat{A}\rangle=\mathrm{Tr}[\hat{A}\hat{\rho}]/\mathrm{Tr}[\hat{\rho}$]. The spin operator in the $j$-direction ($j=x,y,z$) is expressed as $\hat{F}_{j}=\sum_{m,m^{\prime}=-2}^{2}[f_{j}]_{m,m^{\prime}}\hat{a}_{m}^{\dagger}\hat{a}_{m^{\prime}}$, where $f_{j}$ are spin-2 matrices. For the present initial state, $\langle\hat{F}_{j}\rangle$ always vanish as the system evolves. We define the transverse magnetization as
$S_{\perp}=\sqrt{\langle\hat{F}_{x}^{2}+\hat{F}_{y}^{2}\rangle}.$ (43)
To examine whether the state converges to the fully magnetized steady state in Eq. (26) with Eqs. (24) and (25) (or with Eqs. (40) and (41) for $q\neq 0$), we introduce the normalized spin-squared operator as
$\hat{\textbf{F}}^{2}_{\mathrm{norm}}=\frac{\hat{F}_{x}^{2}+\hat{F}_{y}^{2}+\hat{F}_{z}^{2}}{2\hat{N}(2\hat{N}+1)}.$ (44)
To avoid divergence arising from $N=0$, we define the density operator restricted to the $N>0$ subspace as $\hat{\rho}_{N>0}=\oplus_{N=2}^{20}\hat{\rho}_{N}$. The expectation value of Eq. (44) is taken with respect to this restricted state as
$F^{2}_{\mathrm{norm}}=\frac{\mathrm{Tr}[\hat{\textbf{F}}^{2}_{\mathrm{norm}}\hat{\rho}_{N>0}]}{\mathrm{Tr}[\hat{\rho}_{N>0}]}.$ (45)
This quantity becomes unity for the fully magnetized steady state in Eq. (26). Let $P_{N}$ be the probability of finding the system in a state with $N$ particles,
$P_{N}=\frac{\mathrm{Tr}[\hat{\rho}_{N}]}{\mathrm{Tr}[\hat{\rho}]}$. (46)
@@
$Q_{m}=\frac{\mathrm{Tr}[\hat{\mathcal{P}}_{m}\hat{\rho}_{N}]}{\mathrm{Tr}[\hat{\rho}_{N}]},$ (47)
@@
Figure 1: Time evolution for $q=0$ with the initial state in Eq. (42). (a) Average particle number $\langle\hat{N}\rangle$ (blue or dark gray line) and normalized total magnetization $F^{2}_{\mathrm{norm}}$ (red or light gray line). Inset: Log-log plot of the average particle number $\langle\hat{N}\rangle$ (blue or dark gray line) and the expression in Eq. (48) (black line) as a function of the scaled time $T=bt/V^{\mathrm{eff}}+N_{\mathrm{ini}}^{-1}$ with $b=0.0081\times 10^{-11}\mathrm{cm}^{3}/\mathrm{s}$. (b) Particle number distributions $P_{N}$ at (i) $t=0$, (ii) $0.44~\mathrm{s}$, and (iii) $8.8~\mathrm{s}$. The dotted line in (iii) represents a Poisson distribution restricted to even integers in Eq. (49) with $\eta=3.08$. (c) Distributions $Q_{m}$ of the $z$-magnetization at $8.8~\mathrm{s}$. The dotted lines correspond to Eq. (51). See Appendix for $Q_{m}$ in the subspaces with other $N$.